OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). cos x = sin x

ganeshie8 (ganeshie8):

divide both sides with sinx

OpenStudy (anonymous):

then i get cos/sin but how do i relate that to interval 0 and 2pi

OpenStudy (anonymous):

why would you divide both sides by sin(x)

OpenStudy (raden):

square both sides : cos^2 x = sin^2 x cos^2 x = 1 - cos^2 x 2cos^2 x = 1 2cos^2 x - 1 = 0 (sqrt(2)cosx + 1)(sqrt(2)cosx - 1) = 0 now, set each factor be zero then solve for x

OpenStudy (anonymous):

oh thats clever

OpenStudy (anonymous):

cotx = 1 tanx = 1 x = pi/4, + ?

OpenStudy (anonymous):

your a clever little baby RadEn, I assume thats you in your profile picture

OpenStudy (raden):

that's my son's pic heh :)

OpenStudy (anonymous):

so pi/4 and 5pi/4? or is it 7pi/4

OpenStudy (anonymous):

since tanx period is pi, pi+pi/4 also satisfies the solution so, x = pi/4, pi+pi/4 = pi/4, 5pi/4

OpenStudy (anonymous):

Thank you sara!

OpenStudy (anonymous):

yw :)